# ObjectCompose: Evaluating Resilience of Vision-Based Models on Object-to-Background Compositional Changes

### v2

@@ Abstract
Given the large-scale multi-modal training of recent vision-based models and their generalization capabilities, understanding the extent of their robustness is critical for their real-world deployment. In this work, our goal is to evaluate the resilience of current vision-based models against diverse object-to-background context variations. The majority of robustness evaluation methods have introduced synthetic datasets to induce changes to object characteristics (viewpoints, scale, color) or utilized image transformation techniques (adversarial changes, common corruptions) on real images to simulate shifts in distributions. Recent works have explored leveraging large language models and diffusion models to generate changes in the background. However, these methods either lack in offering control over the changes to be made or distort the object semantics, making them unsuitable for the task. Our method, on the other hand, can induce diverse object-to-background changes while preserving the original semantics and appearance of the object. To achieve this goal, we harness the generative capabilities of text-to-image, image-to-text, and image-to-segment models to automatically generate a broad spectrum of object-to-background changes. We induce both natural and adversarial background changes by either modifying the textual prompts or optimizing the latents and textual embedding of text-to-image models. This allows us to quantify the role of background context in understanding the robustness and generalization of deep neural networks. We produce various versions of standard vision datasets (ImageNet, COCO), incorporating either diverse and realistic backgrounds into the images or introducing color, texture, and adversarial changes in the background. We conduct thorough experimentation and provide an in-depth analysis of the robustness of vision-based models against object-to-background context variations across different tasks. Our code is available at https://github.com/Muhammad-Huzaifaa/ObjectCompose
@@
Deep learning-based vision models have achieved significant improvement in diverse vision tasks. However, the performance on static held-out datasets does not capture the diversity of different object background compositions present in the real world. In order for these models to be deployed in security-critical applications, analyzing the robustness of these models under diverse changes in the distribution of the data is crucial. Previous works have shown that vision models are vulnerable to a variety of image alterations, including common corruptions (e.g., snow, fog, blur) [21, 41], domain shifts (e.g., paintings, sketches, cartoons)[19, 20], and changes in viewpoint (e.g., pose, shape, orientation) [5, 27, 3]. Additionally, carefully designed perturbations can be added to images to create adversarial examples that are imperceptible to humans but can fool the decision-making of vision models [58, 16].
@@
Recent works[45, 35] have focused on leveraging existing foundational models to forge new ways to evaluate the resilience of uni/multi-modal vision models. In [45], large language models and text-to-image diffusion models are used for generating diverse semantic changes in real images. However, their method relies on prompt-to-prompt editing[23], allowing only the modification of specific words in the prompt for making changes. Furthermore, due to absence of strong guidance between the object and background during image editing, it often leads to distortion in the object semantics. In [35], diffusion models are used for background editing in real images, and ImageNet-E(diting) dataset is introduced for benchmarking. However, their primary focus is on varying the texture complexity of background by introducing a frequency-based loss for guidance in the denoising process of diffusion model. This imposes limitations on the type of background-compositions that can be achieved. In this work, we address the above limitations and develop a framework for investigating the resilience of vision-based models against diverse object-to-background changes. We accomplish this task by utilising the complementary strengths of current image-to-text[33], image-to-segment[31] and text-to-image[24, 47] foundational models to create a framework that is significantly more adept at handling complex background changes. Our approach preserves the semantics of the original object (Figure 1) by conditioning the text-to-image diffusion model on object boundaries and textual descriptions generated by image-to-segment [31] and image-to-text [33] models. We guide the diffusion model by adding the desired textual description or optimizing its latent visual representation and textual embedding for generating diverse natural and adversarial background changes. Furthermore, we generate datasets covering diverse background changes on a selected subset of ImageNet[10] and COCO validation set [36] enabling comprehensive evaluation of modern uni-modal and multi-modal models across different tasks. Our contribution are as follows:
@@
Our proposed background changes result in an average performance drop of on classifier models compared to the baseline method, and a more substantial drop of when exposed to adversarial background alterations (see Table 1).
@@ 2 Related Work
Adversarial and Counterfactual Manipulations. Researchers have uncovered that subtle, carefully designed alterations to an image, imperceptible to human observers, have the ability to deceive deep learning models [58, 16, 32]. These perturbations, constructed using gradient-based methods, serve as a worst-case analysis in probing the model’s robustness within specified distance norm metrics ( or ). Another strategy entails applying unbounded perturbations to specific image patches, thereby conserving object semantics while inducing model confusion [52, 12]. Recent studies also leverage generative models to create semantic adversarial alterations in images [56, 17, 26, 7, 6, 45, 35]. In [45], LANCE framework is proposed which utilises fine-tuned large language models to get the modified textual prompt for editing of real images to evaluate model’s robustness. However, this framework is not suitable for studying object-to-background compositional changes since the prompt-to-prompt[23] based image editing method leads to global changes in the scene, often altering the object semantics. This necessitates hyper-parameter tuning of parameters used for prompt-to-prompt editing, leading to generation of multiple edited image versions and selecting the one most faithful to the original image. In [35], using already available masks of ImageNet dataset [13], diffusion model is utilised to alter the background of images by varying its texture. A complexity loss based on gray-level co-occurence matrix[18] of the image is used during the denoising process to vary the complexity of the background. On the contrary, our method can induce natural background variations through textual guidance and optimize the latent space of the diffusion model for adversarial changes.
@@ 3 Method
We introduce ObjectCompose, our method for generating diverse language-guided object-to-background compositional changes to evaluate the resilience of current vision-based models. ObjectCompose does this by exploiting the complementary strength of foundational models; image-to-segment, and image-to-text to guide object-preserving diffusion inference for natural and adversarial background-to-object context variations (see Figure 2). Our automated approach effectively generates different datasets under varying distribution shifts, which can be used for benchmarking vision and vision language models.
In Section 3.1, we provide the preliminaries of text-to-image, image-to-text, and image-to-segment models that are used in our method. Subsequently, in Section 3.2, we explain the working of our method in detail.
@@ 3.1 Preliminaries
Diffusion Models. Diffusion models have made remarkable advancements in generating high quality images and refining them based on textual guidance. During training, diverse noisy versions of the clean image are fed into the diffusion model at different time steps . The objective for the model is to learn the specific noise added at each time step. The training process of diffusion models comprises of two stages; in the forward process (first stage) gaussian noise sampled from a normal distribution is gradually added to image according to a variance schedule . Using the reparameterization trick, we can get the noisy image at any time step as follows:
Here, and . As , , which implies , all the information of the original image will be lost. Typically, diffusion models are conditioned on the time step and additional factors like the class label or textual description . However, recent works have extended this conditioning to also consider the clean image and its corresponding mask, enabling specialized image editing tasks [47, 49]. Next, in the reverse process (second stage) based on the conditioning, a model is learned to approximate the gaussian parameters at each time step for the true reverse conditional distribution. Essentially, the training objective minimizes the error between the estimated and actual noise added to the image at different time steps :
where is the embedding of conditional guidance through either class label or caption of the image and is any additional conditioning, such as mask or layout of the scene.
Foundational Models. BLIP-2 [33] presents an efficient vision-language pre-training approach that utilizes a lightweight Querying Transformer (QFormer) to bridge the modality gap between pre-trained vision and large language models (LLMs). This framework initially passes the image through a pre-trained vision encoder and extracts relevant features via the QFormer. This information is then passed to the pre-trained LLM to obtain a descriptive caption of the image.
Recently, Segment Anything Model (SAM)[31], a image segmentation model was introduced,that undergoes pre-training on an extensive dataset of high quality images. SAM employs prompts, which can manifest in various forms such as point sets, boxes, masks, or textual input, to demarcate objects within an image. The image initially undergoes an encoding process through a large transformer-based image encoder. Subsequently, both the features extracted from the image and the embeddings of the prompt from a prompt encoder traverse a lightweight decoder, yielding the desired segmentation mask.
@@ 4 Experimental Protocols
Dataset Preparation. For classification, we initially gathered 30k images from the ImageNet validation set [10], which are correctly classified with high success rate using an ensemble of models; ViT-T, ViT-S [11], Res-50, Res-152 [19], Dense-161 [25], Swin-T, and Swin-S [37]. In order to create a high-quality dataset for our object-to-context variation task, we remove image samples where the boundary between foreground and background is not distinct, e.g., "mountain tent" where the mountain might appear in the background of the tent. This processing results in 15k images. Then for foreground semantic preservation, we utilize a compute-efficient variant of SAM, known as FastSAM [64] with class labels as prompts to generate segmentation masks of the foreground object. However, FastSAM encounter challenges in accurately segmenting objects in all images. To address this, we selected images where the mask-creation process demonstrated exceptional accuracy and generated a clear separation between the object of interest and its background. This meticulous selection process yield a curated dataset comprising 5,505 images, representing a subset of 582 ImageNet classes. We refer to this dataset as ImageNet-B. More details are presented in Appendix 0.A.17. Due to the computational cost involved in adversarial background optimization and running baseline methods, we select a subset of 1000 images from 500 classes of ImageNet-B by sampling two images from each class for comparison. We refer to this dataset as . Rest of our experiments are performed on the full ImageNet-B dataset.
For object detection, we carefully filtered 1,127 images manually from the COCO 2017 validation set [36], with a clear distinction between foreground objects and their background. We refer to these images as COCO-DC. We use COCO-DC to evaluate both detection and classification. The dataset can have multiple objects in the image. To use this dataset for classification we train the above-mentioned models on the COCO train dataset using labels associated with the object that occupies the highest mask region in the image and evaluated on our generated dataset.
Diffusion Parameters. We use the pre-trained Inpaint Stable Diffusion v2[47] as our text-to-image model and set the guidance parameter to , and use the DDIM sampling [55] with timesteps. We craft adversarial examples on using Res-50[19] as the classifier model and maximize the adversarial loss shown in Eq.6 for iterations. For COCO-DC, we maximize the loss in the feature space of the model. Both the text embedding of the prompt (initialized with ) and denoised latent are optimized from denoising time step using AdamW [39] with a learning rate of . All experiments were conducted using a single NVIDIA-A100 GPU.
@@
Evaluation Metrics: We report results using the top-1 accuracy (%), Intersection Over Union (IoU), Average Precision(AP) and Recall(AR), and CLIP text similarity score for classification, segmentation, object detection, and captioning tasks, respectively.
Background Conditioning. During inducing background variations, we use the following text prompts templates; Class Label: "A picture of a class" where class is the class name of the image, Caption: "captions from BLIP-2 model, Color: "A picture of ___ background" where __ is replaced with red, green, blue, and colorful, Texture: "A picture of __ background where __ is replaced with textured, intricately rich textures, colorful textures, distorted textures, Adversarial: "captions from BLIP-2 model". Note that for adversarial setting the text prompts get updated after optimization. In the case of the ImageNet-E method[35], default values of are employed to regulate the strength of texture complexity introduced in both natural and adversarial background changes. Likewise, for LANCE[45], the default prompt specified in the paper is utilized, passing it to the large language model to generate variations of captions that reflect background changes. We observe similar trends across different color and texture prompts and report the worst-performing one. Detailed analysis across different color and texture prompts is provided in Appendix 0.A.5.
@@ 4.1 Comparison with Baseline Methods
| Class label | 78.49 | 83.66 | 81.58 | 76.60 | 77.00 | 82.09 | 84.50 | 80.55 |
| BLIP-2 Captions | 68.79 | 72.29 | 71.49 | 65.20 | 68.40 | 71.20 | 75.40 | 70.40 |
@@
Object Semantic distortion: It’s noteworthy to mention that in both Table 1 and 2, we observe a significant drop in performance of models across background changes induced by LANCE method. However, we discover that the drop in performance is not necessarily due to the induced background changes, rather than distorting the object semantics, making it unsuitable for evaluating object-to-background context. This observation is supported by Figure 5, where we evaluate performance on original and LANCE generate images while masking the background. A significant performance drop is evident across all models, emphasizing the distortion of object semantics. For a more detailed discussion and visualizations, please refer to Appendix 0.A.3 and 0.A.4.
@@ 4.2 Further Evaluations
Segmentation and Detection. We observe a consistent decrease in AP scores on the task of object detection and instance segmentation across different background variations generated on COCO-DC(see Table 5). The adversarial background results in the lowest AP scores, but still remains at a reasonable level given that the adversarial examples are generated using a classification model, with limited cross-task transferability. Moreover, our qualitative observations suggest detection and segmentation models exhibit greater resilience to changes in the background compared to classifiers (see Figure 8 and Appendix 0.A.10).
5 Conclusion
In this study, we propose ObjectCompose, a method for generating object-to-background compositional changes. Our method addresses the limitations of current works, specifically distortion of object semantics and diversity in background changes. We accomplish this by utilizing the capabilities of image-to-text and image-to-segmentation foundational models to preserve the object semantics, while we optimize for diverse object-to-background compositional changes by modifying the textual prompts or optimizing the latents of the text-to-image model. ObjectCompose offers a complimentary evaluation protocol to the existing ones, for comprehensive evaluations across current vision-based models to reveal their vulnerability to background alterations. We anticipate that our insights will pave the way for a more thorough evaluation of vision models, consequently driving the development of more effective methods for improving their resilience.

### v4

@@ Abstract
Given the large-scale multi-modal training of recent vision-based models and their generalization capabilities, understanding the extent of their robustness is critical for their real-world deployment. In this work, our goal is to evaluate the resilience of current vision-based models against diverse object-to-background context variations. The majority of robustness evaluation methods have introduced synthetic datasets to induce changes to object characteristics (viewpoints, scale, color) or utilized image transformation techniques (adversarial changes, common corruptions) on real images to simulate shifts in distributions. Recent works have explored leveraging large language models and diffusion models to generate changes in the background. However, these methods either lack in offering control over the changes to be made or distort the object semantics, making them unsuitable for the task. Our method, on the other hand, can induce diverse object-to-background changes while preserving the original semantics and appearance of the object. To achieve this goal, we harness the generative capabilities of text-to-image, image-to-text, and image-to-segment models to automatically generate a broad spectrum of object-to-background changes. We induce both natural and adversarial background changes by either modifying the textual prompts or optimizing the latents and textual embedding of text-to-image models. This allows us to quantify the role of background context in understanding the robustness and generalization of deep neural networks. We produce various versions of standard vision datasets (ImageNet, COCO), incorporating either diverse and realistic backgrounds into the images or introducing color, texture, and adversarial changes in the background. We conduct thorough experimentation and provide an in-depth analysis of the robustness of vision-based models against object-to-background context variations across different tasks. Our code and evaluation benchmark will be available at https://github.com/Muhammad-Huzaifaa/ObjectCompose.
@@
Deep learning-based vision models have achieved significant improvement in diverse vision tasks. However, the performance on static held-out datasets does not capture the diversity of different object background compositions present in the real world. Previous works have shown that vision models are vulnerable to a variety of image alterations, including common corruptions (e.g., snow, fog, blur) [20, 40], domain shifts (e.g., paintings, sketches, cartoons)[18, 19], and changes in viewpoint (e.g., pose, shape, orientation) [5, 26, 3]. Additionally, carefully designed perturbations can be added to images to create adversarial examples that are imperceptible to humans but can fool the decision-making of vision models [56, 16].
@@
The main motivation of our work is to understand how object-to-background compositional changes in the scene impact uni/multi-modal model performance. Recent works[44, 34] have focused on leveraging existing foundational models to forge new ways to evaluate the resilience of uni/multi-modal vision models. In [44], large language models and text-to-image diffusion models are used for generating diverse semantic changes in real images. However, their method employs the prompt-to-prompt method[22] for image editing, allowing limited word changes in the textual prompt to preserve object semantics. We also observe that it suffers from object distortion due to absence of strong guidance between the object and background during image editing. In [34], diffusion models are used for background editing in real images, and ImageNet-E(diting) dataset is introduced for benchmarking. However, their use of a frequency-based loss for guiding the generation process of diffusion models limit the control to attribute changes in the background. This imposes limitations on the type of background-compositions that can be achieved.
In this work, we develop a framework to investigate the resilience of vision models to diverse object-to-background changes. Leveraging the complementary strengths of image-to-text [32], image-to-segment [30], and text-to-image [24, 46] models, our approach better handles complex background variations. We preserve object semantics (Figure 1) by conditioning the text-to-image diffusion model on object boundaries and textual descriptions from image-to-segment and image-to-text models. We guide the diffusion model by adding the desired textual description or optimizing its latent visual representation and textual embedding for generating diverse natural and adversarial background changes. Additionally, we produce datasets with varied backgrounds from subsets of ImageNet [10] and COCO [35], facilitating the evaluation of uni-modal and multi-modal models. Our contributions are as follows:
@@
Our proposed background changes yield an average performance drop of on classifier models compared to the baseline method, and a substantial drop of when exposed to adversarial changes (see Table 1).
@@ 2 Related Work
Adversarial and Counterfactual Manipulations. Researchers have uncovered that subtle, carefully designed alterations to an image, imperceptible to human observers, have the ability to deceive deep learning models [56, 16, 31]. These perturbations, constructed using gradient-based methods, serve as a worst-case analysis in probing the model’s robustness within specified distance norm metrics ( or ). Another strategy entails applying unbounded perturbations to specific image patches, thereby conserving object semantics while inducing model confusion [51, 12]. Several studies leverage generative models to create adversarial alterations in images. [7, 6] utilise diffusion model and GANs to introduce global adversarial perturbations in the image with strong constraint to semantic changes in order to preserve the original layout of the scene. More recent works[34, 44] are more closely related to our goal of evaluating vision-based models on object-to-background compositional changes in the scene. In [44], LANCE framework is proposed which utilises fine-tuned large language models to get the modified textual prompt for editing of attributes in real images. However, this framework is not ideal for studying object-to-background compositional changes since the prompt-to-prompt[22] based image editing method often leads to global changes in the scene, often altering the object semantics. This necessitates hyper-parameter tuning of parameters used for prompt-to-prompt editing, leading to generation of multiple edited image versions and selecting the one most faithful to the original image. In [34], using already available masks of ImageNet dataset[13], diffusion model is utilised to alter the background of images by varying its texture. A complexity loss based on gray-level co-occurence matrix[17] of the image is used during the denoising process to vary the complexity of the background. A concurrent work[61] evaluates the resilience of models on synthetic images where both the object and background are generated using a diffusion model. In contrast to previous works, our method induces natural/adversarial background variations in real images through textual guidance and optimization of latent space of the diffusion model, all while maintaining the integrity of the object semantics. Our method can be applied to standard vision datasets to generate diverse background variations, providing a robust benchmark for evaluating vision-based models.
@@ 3 Method
We introduce ObjectCompose, a method for generating diverse language-guided object-to-background compositional changes to evaluate the resilience of vision models. ObjectCompose leverages the complementary strengths of image-to-segment and image-to-text models to guide object-preserving diffusion for natural and adversarial background variations (Figure 2). Our automated approach generates datasets under varying distribution shifts, useful for benchmarking vision and vision-language models.
In Section 3.1, we outline the preliminaries of the foundational models used. In Section 3.2, we detail our method.
@@ 3.1 Preliminaries
Diffusion Models. Diffusion models have significantly advanced in generating high-quality images and refining them based on textual guidance. During training, noisy versions of the clean image are input to the model at various time steps , with the goal of learning the noise added at each step. Training consists of two stages: in the forward process (first stage), Gaussian noise from a normal distribution is incrementally added to according to a variance schedule . Using reparameterization, the noisy image at any time step is:
Here, and . As , , meaning and all information from is lost. Diffusion models are typically conditioned on , class label , or textual description , with recent extensions incorporating image and its mask for image editing tasks [46, 48].
In the reverse process (second stage), the model learns to approximate the Gaussian parameters at each time step for the reverse conditional distribution. The objective minimizes the error between the predicted and actual noise at each time step:
where is the embedding of the conditional guidance, and represents any additional conditioning, such as masks or scene layouts.
Foundational Models. BLIP-2 [32] introduces an efficient vision-language pre-training approach using a lightweight Querying Transformer (QFormer) to bridge the gap between pre-trained vision and large language models (LLMs). Images are processed by a vision encoder, with relevant features extracted via QFormer and passed to the LLM to generate descriptive captions.
Recently, Segment Anything Model (SAM)[30], an image segmentation model, was introduced that undergoes pre-training on an extensive dataset of high-quality images. SAM uses prompts—such as points, boxes, masks, or text—to identify objects in images. The image is encoded by a transformer-based encoder, and the extracted features, combined with prompt embeddings, are processed by a lightweight decoder to produce the segmentation mask.
@@ 4 Experimental Protocols
Dataset Preparation. For classification, we initially gathered 30k images from the ImageNet validation set [10], which are correctly classified with high success rate using an ensemble of models; ViT-T, ViT-S [11], Res-50, Res-152 [18], Dense-161 [25], Swin-T, and Swin-S [36]. In order to create a high-quality dataset for our object-to-context variation task, we remove image samples where the boundary between foreground and background is not distinct, e.g., "mountain tent" where the mountain might appear in the background of the tent. This processing results in 15k images. Then for foreground semantic preservation, we utilize a compute-efficient variant of SAM, known as FastSAM [63] with class labels as prompts to generate segmentation masks of the foreground object. However, FastSAM encounter challenges in accurately segmenting objects in all images. To address this, we selected images where the mask-creation process demonstrated exceptional accuracy and generated a clear separation between the object of interest and its background. This meticulous selection process yield a curated dataset comprising 5,505 images, representing a subset of 582 ImageNet classes. We refer to this dataset as ImageNet-B. Due to the computational cost involved in adversarial background optimization and running baseline methods, we select a subset of 1000 images from 500 classes of ImageNet-B by sampling two images from each class for comparison. We refer to this dataset as . Rest of our experiments are performed on the full ImageNet-B dataset.
For object detection, we manually filtered 1,127 images from the COCO 2017 validation set [35], ensuring a clear distinction between foreground objects and background, referred to as COCO-DC. This dataset, containing multiple objects per image, is used for both detection and classification. For classification, models are trained on the COCO train dataset using the label of the object with the largest mask region and evaluated on our generated dataset. Additional details and dataset comparisons are provided in Appendix 0.A.16.
Diffusion Parameters. We use the pre-trained Inpaint Stable Diffusion v2[46] as our text-to-image model and set the guidance parameter to , and use the DDIM sampling [54] with timesteps. We craft adversarial examples on using Res-50[18] as the classifier model and maximize the adversarial loss shown in Eq.6 for iterations. For COCO-DC, we maximize the loss in the feature space of the model. Both the text embedding of the prompt (initialized with ) and denoised latent are optimized from denoising time step using AdamW [38] with a learning rate of .
@@
Evaluation Metrics: We use the top-1 accuracy (%), Intersection Over Union (IoU), Average Precision(AP) and Recall(AR), and CLIP text similarity score for classification, segmentation, object detection, and captioning tasks, respectively.
Background Conditioning. To induce background variations, we use the following text prompt templates: Class Label: "A picture of a class" where class is the image’s class name; Caption: "captions from BLIP-2"; Color: "A picture of ___ background" where ___ is red, green, blue, or colorful; Texture: "A picture of __ background" with ___ replaced by textured, rich textures, colorful textures, distorted textures; Adversarial: "captions from BLIP-2" with updated prompts after optimization. For ImageNet-E[34], default values of are employed to regulate the strength of texture complexity. For LANCE[44], we use the default prompt to generate background variations via a large language model. We report the worst-performing prompt across colors and textures, with detailed analysis in Appendix 0.A.5.
@@ 4.1 Comparison with Baseline Methods
| Class label | 78.49 | 83.69 | 88.79 | 76.60 | 77.00 | 82.09 | 84.50 | 81.59 |
| BLIP-2 Captions | 68.79 | 72.29 | 79.19 | 65.20 | 68.40 | 71.20 | 75.40 | 71.49 |
@@
Object Semantic distortion: It’s noteworthy to mention that in both Table 1 and 2, we observe a significant drop in performance of models across background changes induced by LANCE method. However, we discover that the drop in performance is not necessarily due to the induced background changes, rather than distorting the object semantics, making it unsuitable for evaluating object-to-background context. This observation is supported by Figure 5, where we evaluate performance on original and LANCE generate images while masking the background. A significant performance drop is evident across all models, emphasizing the distortion of object semantics. We provide a detailed discussion with FID[23] comparison and visualizations in Appendix 0.A.3 and 0.A.4.
@@ 4.2 Further Evaluations
Segmentation and Detection. We observe a consistent decrease in AP scores on object detection and instance segmentation tasks across background variations generated on COCO-DC(see Table 5). The adversarial background results in the lowest AP scores, but still remains at a reasonable level given that the adversarial examples are generated using a classification model, with limited cross-task transferability. Moreover, our qualitative observations suggest detection and segmentation models exhibit greater resilience to changes in the background compared to classifiers (see Figure 8 and Appendix 0.A.10).
Image Captioning. Table 5 shows the CLIP scores between captions from clean and generated images using the BLIP-2 model. Scores decrease with color, texture, and adversarial background changes (Appendix 0.A.9 for qualitative results).
@@
5 Conclusion
In this study, we propose ObjectCompose, a method for generating object-to-background compositional changes. Our method addresses the limitations of current works, specifically distortion of object semantics and diversity in background changes. We accomplish this by utilizing the capabilities of image-to-text and image-to-segmentation foundational models to preserve the object semantics, while we optimize for diverse object-to-background compositional changes by modifying the textual prompts or optimizing the latents of the text-to-image model. ObjectCompose offers a complimentary evaluation protocol to the existing ones, for comprehensive evaluations across current vision-based models to reveal their vulnerability to background alterations. In Appendix 0.A.18, we elaborate on the initial insights gained from our work and discuss current limitations and future directions.